# Imaginarity of quantum channels: Refinement and Alternative

### v1

@@ Abstract
At present, quantum channels have been widely concerned, and many ways to quantify quantum channels have been proposed, which has led to the generation of many resource theories for quantum channels. We add strong monotonicity and convexity to the requirement of imaginarity measure of quantum channels to make the measure proper. We also introduce an alternative framework to simplify the process of verifying whether a quantifier is a proper measure. We present three imaginarity measures of quantum channels via on the robustness, the trace norm and entropy, respectively. Some properties are also given.
Keywords: quantum resource theories, quantum channels, imaginarity measures, alternative framework
@@ 1 Introduction
Imaginary numbers provide a broad perspective and additional methods for modern scientific research. Quantum resource theory can establish new connections among various research directions. With the development of quantum resource theory, the imaginarity resource theory [1] proposed in recent years has attracted increasing attention. Some examples of applications, the imaginarity plays a crucial role in state discrimination [2], the imaginarity operation is more advantageous to distinguish the quantum channel without the help of the auxiliary system [3]. Many imaginarity measures have been found for quantum states, the fidelity [2], the convex roof [4], the geometric [3] and some imaginarity measures induced by entropies [5, 6], some of which can establish relations due to their own mathematical properties [6, 7].
Just like quantum states, quantum channels also deserve attention. Some studies begin to investigate the properties of quantum channels [8, 9, 10, 11, 12]. Resource theories of quantum channels was established in [13], and many resource theories of quantum channels have been proposed, such as coherence of quantum channels [14], entanglement of quantum channels [15] and so on [16, 17]. Based on this, more measures for quantum channels have been found [18, 19, 20, 21].
In order to simplify the process of verifying whether a quantifier can be a coherence measure, alternative framework for quantifying coherence have been proposed [22] in the coherence resource theory. Similar alternative frameworks can be established in many other resource theories, such as imaginarity resource theory [7], coherence of quantum channels [23].
The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been visited again recently [25]. It can be seen that the concept is worthing considing. Similar to the coherence of quantum channels, in this paper, we refine the imaginarity of quantum channels and introduce alternative framework to facilitate the verification of whether a quantifier can be a suitable measure. We also show three imaginarity measures of quantum channels. Some properties of the measure are discussed.
2 Refinement for the imaginarity of quantum channels
@@
$J_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)=\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
@@
The Choi state of $\phi$ is $\frac{J_{\phi}}{|A|}\triangleq\chi_{\phi}$. We call channel $\phi$ a real channel if $\chi_{\phi}$ is an real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels. The free channel is defined as the real channel. This is exactlt the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$. It can be known that [14, 26] superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where $\tilde{\Theta}$ is the channel corresponding to the superchannel. We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
@@
In [24], the author presented a framework for quantifying the imaginarity of quantum channels, they introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
@@
Strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them. So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes the more familiar and complete.
@@
Monotonicity under composition with any real channels, that is,
$(a)$. Left composition: for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$, we have: $C(\psi\circ\phi)\leqslant C(\phi)$.
$(b)$. Right composition: for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$, we have: $C(\phi\circ\psi)\leqslant C(\phi)$.
@@
We only prove $(a)$, because $(b)$ is analogous.
For $\psi\in\mathcal{RC}_{AB}$ and $\phi\in\mathcal{C}_{BD}$, we consider a superchannel $\Theta_{\psi}$ with $\Theta_{\psi}=\psi\circ\phi$. Due to the condition $\bm{(C_{2})}$, we try to explain $\Theta\in\mathcal{RSC}_{ABAD}$. In fact, when superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate imaginarity resource since the channel $\psi$ is real. We can know that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition. ∎
We can also see that if $\psi$ is invertible, $C(\psi\circ\phi)=C(\phi)$ holds.
@@
The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure .
@@
It is obvious that $C(\phi)\geqslant 0$ and due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, certainly it indicate $\phi\in\mathcal{RC}$. For any real superchannel $\Theta$, we have $C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}C_{j}(\Theta(\phi))=C(\Theta(\phi))$.
@@
If $C$ is an imaginarity measure for channel $\phi$, then, quantifier $\hat{C}$ as:
@@
is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with $q_{j}\geqslant 0$.
@@
We can notice that, for any $\left\{q_{j}\right\}$, let every $\phi_{j}=\phi$, we can get $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$. For any quantum channel $\phi$, we take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we denote $a_{k,m}=\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$, then
$\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$\geqslant$ $\sum_{m}p_{m}\sum_{k}C(\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})\geqslant\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
The first inequality is due to the strong monotonicity of $C$, the second inequality is due to the convexity of $C$ and $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$.
To verify convexity of $\hat{C}$ for simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$. We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the equations hold:
@@
For ease of writing, if $m\neq n$, we supplement $\left\{0,\bm{0}\right\}$ for the less set so that both sets have the same quantity, that is, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$. Substitute the new symbol $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ for $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively. We have
$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})\geqslant\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})=\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
In this way, it can be known that $\hat{C}$ is a proper imaginarity measure. ∎
@@ 3 Alternative framework for the imaginarity of quantum channels
In the previous section we refined the framework. However, the condition $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are difficult to verify in some cases. Here, inspired by the similar alternative framework in many resource theory [22, 23], we have the following result.
@@
$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds. The condition $\bm{(C_{5})}$ is:
$\bm{(C_{5})}$. Additivity: For any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$, then $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ holds.
@@
The proof is long and we divide it into steps.
$\bm{(C_{3})}+\bm{(C_{4})}\Rightarrow\bm{(C_{5})}$:
(STEP 1).
@@
$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
@@
Due to $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$, we can know that $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
@@
$\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
@@
It’s obvious that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$. Let $\chi_{\tilde{\phi}_{1}}=Q_{1}\chi_{\phi}Q_{1}^{\dagger}/p_{1}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|$. Similarly we obtain $\chi_{\tilde{\phi}_{2}}$. At this point, the requirement of $\bm{(C_{3})}$ has been satisfied, that is
$\begin{cases}\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}\text{ with }\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger},\\ p_{m}=\text{Tr}(Q_{m}\chi_{\phi}Q_{m}^{\dagger}),\chi_{\tilde{\phi}_{m}}=Q_{m}\chi_{\phi}Q_{m}^{\dagger}/p_{m},\end{cases}$
we have
$C(\chi_{\phi})\geqslant p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}}).$ (1)
(STEP 2).
@@
then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, we have
$\phi(\rho)$ $=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)=\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$

### v2

@@ Abstract
In this paper, we introduce the framework for quantifying the imaginarity of quantum channels. Besides, an alternative framework is given together to simplify the process of verifying the condition. We present three imaginarity measures of quantum channels based on the robustness, the trace norm, and the entropy, respectively. Some properties are also given.
keywords:
quantum resource theories , quantum channels , imaginarity measures , alternative framework
[…]
organization=School of Mathematics, addressline=Harbin Institute of Technology, city=Harbin, postcode=150001, country=China
@@ 1 Introduction
Imaginary numbers offer a broad perspective and provide additional methodologies for modern scientific research. Quantum resource theory facilitates the establishment of new connections across various research directions. In recent years, the concept of imaginarity resource theory [1], which has emerged alongside the development of quantum resource theory, has garnered increasing attention. Several applications can illustrate its significance. For instance, imaginarity plays a critical role in state discrimination [2]. The imaginarity operation proves to be particularly advantageous for distinguishing quantum channels without the need for an auxiliary system [3]. Numerous imaginarity measures have been identified for quantum states, including those based on fidelity [2], the convex roof [4], geometric considerations [3], and some induced by entropies [5, 6]. Some of them establish relations due to their mathematical properties [6, 7]. Just as quantum states warrant attention, quantum channels also merit investigation. Recent studies have begun to explore the properties of quantum channels [8, 9, 10, 11, 12]. The resource theories of quantum channels were first established in [13], with several subsequent theories proposed, such as coherence in quantum channels [14], entanglement in quantum channels [15], and more [16, 17]. Building upon these frameworks, additional measures for quantum channels have been introduced [18, 19, 20, 21].
The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been revisited recently [25]. This indicates that the concept is of significant interest and worth considering. The structure of this paper is as follows: Similar to the coherence of quantum channels, we refine the concept of imaginarity for quantum channels and present an alternative framework that facilitates the verification of whether a given quantifier can serve as a suitable measure. We discuss the properties of imaginarity measures of quantum channels. Additionally, we propose three measures of the imaginarity of quantum channels based on the robustness, the trace norm, and the entropy, respectively.
2 Framework for quantifying the imaginarity of quantum channels
@@
$J_{\phi}$ $=$ $\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)$
$=$ $\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
@@
The Choi state of $\phi$ is $\frac{J_{\phi}}{|A|}\triangleq\chi_{\phi}$. We call channel $\phi$ a real channel if $\chi_{\phi}$ is a real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels. The free channel is defined as the real channel. This is exactly the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels $\Theta$ from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$. The superchannel $\Theta$ can be represented by Choi matrix [14],
$J_{\Theta}$ $=$ $\sum_{j,k,\alpha,\beta}|j\alpha\rangle\langle k\beta|\otimes\Theta(|j\alpha\rangle\langle k\beta|)$
$=$ $\sum_{j,k,\alpha,\beta}\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}|j\alpha j^{\prime}\alpha\rangle\langle k\beta k^{\prime}\beta^{\prime}|,$
where $\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}=\langle j^{\prime}\alpha^{\prime}|\Theta(|j\alpha\rangle\langle k\beta|)|k^{\prime}\beta^{\prime}\rangle$ with $|j\alpha\rangle$ means the tensor of the basis from systems $A$ and $B$.
Similar to the case of channel, superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where
$M_{m}=\sum_{j,j^{\prime},\alpha,\alpha^{\prime}}M_{m,j^{\prime},j,\alpha^{\prime},\alpha}|j^{\prime}\alpha\rangle\langle j\alpha|$
with $\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}=\sum_{m}M_{m,j^{\prime},j,\alpha^{\prime},\alpha}M_{m,k^{\prime},k,\beta^{\prime},\beta}$.
Here, $\tilde{\Theta}$ means the channel corresponding to the superchannel [14, 26]. We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
@@
In [24], the author presented a framework for quantifying the imaginarity of quantum channels. They introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
@@
Actually, strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them. So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes more familiar and complete.
@@
Monotonicity under composition with any real channels:
$(a)$. Left composition: $C(\psi\circ\phi)\leqslant C(\phi)$ for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$.
$(b)$. Right composition: $C(\phi\circ\psi)\leqslant C(\phi)$ for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$.
$(C)$. Right composition: $C(\phi\circ\psi)=C(\phi)$ for any invertible $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$.
@@
We only prove $(a)$, because $(b)$ is analogous. $(c)$ can be obtained from $(a)$ and $(b)$.
For $\psi\in\mathcal{RC}_{AB}$ and $\phi\in\mathcal{C}_{BD}$, we consider a superchannel $\Theta_{\psi}$ with $\Theta_{\psi}=\psi\circ\phi$. When superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate the imaginarity resource since the channel $\psi$ is real. It can be seen that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition. The result holds due to the condition $\bm{(C_{2})}$. ∎
@@
The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any probability distribution $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure.
@@
It is obvious that $C(\phi)\geqslant 0$. Due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, and it indicate $\phi\in\mathcal{RC}$. For any real superchannel $\Theta$, we have $C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}C_{j}(\Theta(\phi))=C(\Theta(\phi))$.
@@
If $C$ is an imaginarity measure for channel $\phi$, then the quantifier $\hat{C}$ as
@@
is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with probability distribution $q_{j}\geqslant 0$.
@@
For any probability distribution $\left\{q_{j}\right\}$, if we suppose every $\phi_{j}=\phi$, we can have $\sum_{j}q_{j}C(\phi_{j})=C(\phi)$. Thus, it follows that $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$. For any quantum channel $\phi$, we take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we denote $a_{k,m}=\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$, then
$\hat{C}(\phi)$
$=$ $\sum_{k}q_{k}C(\phi_{k})$
$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$=$ $\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$\geqslant$ $\sum_{m}p_{m}C(\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$=$ $\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})$
$\geqslant$ $\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
The first inequality follows from the strong monotonicity of $C$. The second inequality arises from the convexity of $C$ and the fact that $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$. The third inequality results from $\hat{C}(\phi)\leqslant C(\phi)$.
To verify the convexity of $\hat{C}$, for simplicity and without loss of generality, we consider only two quantum channels $\phi_{1}$ and $\phi_{2}$. We define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$. We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the following equations hold:
@@
For ease of writing, if $m\neq n$, we supplement the elements $\left\{0,\bm{0}\right\}$ to the smaller set so that both sets have the same number of elements. This gives $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$. Next, we introduce the new symbols $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ to represent $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively. Then we have
$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})$
$=$ $\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})$
$\geqslant$ $\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})$
$=$ $\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
In this way, we confirm that $\hat{C}$ is a valid imaginarity measure. ∎
@@ 3 Alternative framework for the imaginarity of quantum channels
In Section 2, we introduce the framework. However, conditions $\bm{(C_{3})}$ and $\bm{(C_{4})}$ can sometimes be difficult to verify. In coherence resource theory, to simplify the process of verifying whether a quantifier qualifies as a coherence measure, an alternative framework for quantifying coherence has been proposed [22]. Similar alternative frameworks can also be established in many other resource theories, such as imaginarity resource theory [7], and coherence of quantum channels [23]. Inspired by [23], we Introduce the following condition:
$\bm{(C_{5})}$. Additivity: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, where $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$.
When conditions $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold, the conditions $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are equivalent to the condition $\bm{(C_{5})}$. To prove this, we first state the following theorem:
@@
$\bm{(C_{3})}$ and $\bm{(C_{4})}$ give rise to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
@@
STEP 1, obtain: $\bm{C(\phi)\geqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})}$.
@@
$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}$
$=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}$
$=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
@@
Given that $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$. It follows that $\Theta\in\mathcal{RSC}_{ABAB}$.
@@
$\chi_{\phi}$
$=$ $\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$
$=$ $\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
@@
It is clear that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$. Let $\chi_{\tilde{\phi}_{1}}=Q_{1}\chi_{\phi}Q_{1}^{\dagger}/p_{1}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|$. Similarly we can obtain $\chi_{\tilde{\phi}_{2}}$. From the condition $\bm{(C_{3})}$, we have:
$C(\phi)=C(\chi_{\phi})$$\geqslant$ $p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}})$ (1)
$=$ $p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$
STEP 2, obtain: $\bm{C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})}$.
@@
then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, for any $\rho$ we have
$\phi(\rho)$
$=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$
$=$ $\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}$
$+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$